OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. S(1/2, 1), T(1/2, 4)

OpenStudy (shayaan_mustafa):

use two points formula Do you know?

OpenStudy (shayaan_mustafa):

\[(x-x _{1})/(x _{2}-x _{1}) = (y-y _{1})/(y _{2}-y _{1})\]

OpenStudy (anonymous):

i did but i guess i'm not doing it right

OpenStudy (shayaan_mustafa):

send me your step by step answer.

OpenStudy (anonymous):

x=1/2

OpenStudy (shayaan_mustafa):

and y?

OpenStudy (anonymous):

i only need to find x

OpenStudy (anonymous):

since the points being 1/2,1 and 1/2,4 this shows that for the line passing through these point value of x is constan =1/2